OpenStudy (anonymous):

write the equation of a line in slope-intercept fo that is parallel to y=5x-10 and passes throught the point (9,-3) OpenStudy

OpenStudy (anonymous):

(9,-3); y=5x-10. ---------- You would use point-slope equation which is y-y1=m(x-x1). Your slope will stay the same since it is parallel, in which you will plug in 5 for m (m=slope). y-y1=5(x-x1). Next plug in the coordinate (9,-3). The 9 will replace x1 and the -3 will replace y1. y-(-3)=5(x-9) Next simpliy and solve for y. y-(-3)=5(x-9) y+3=5x-45 -3 -3 y=5x-48. Your answer will be y=5x-48.

OpenStudy (anonymous):

if you were to be asked for the PERPENDICULAR rather than parallel, you would take the opposite of the slope given and flip it (the reciprocal)